Official Program of Annual International Workshop on Nonlinear Analysis and Convex Analysis

KYOTO UNIVERSITY

Oiwake-chou Kita-Shirakawa, Sakyou-ku, Kyoto-shi

This Workshop is officially approved by Kyoto University and financially supported by Research Institute for Mathematical Sciences attached to Kyoto University.

Directed by Wataru Takahashi(Tokyo Institute of Technology) and organaized by Shigeo Akashi(Tokyo University of Science)

 Date: The workshop begins at 9:00 on the 3rd of September and it ends at 15:10 on the 5th.
 Place: This workshop is held at Room 115 located on the grand floor

of Research Institute of Mathematical Sciences attached to Kyoto University.

As for the conference venue, please refer to www.kurims.kyoto-u.ac.jp/en/access-01.html.

September 3(Monday)

9:00〜9:30 Hang-Chin Lai (Chung Yuan Christian University) Minimax Programming in Complex Space

9:30〜10:00 Wataru Takahashi (Tokyo Institute of Technology) Nonlinear Ergodic Theorems for Nonexpansive Semigroups in Banach Spaces

10:00〜10:30 Syuuji Yamada, Tamaki Tanaka (Niigata University) and Tetsuzo Tanino (Osaka University) A Successive Approximation Method for Solving Lipschitz Optimization Problems

10:30〜10:40 Tea Break 10:40〜11:10 Do Sang Kim, Hyo Jung Lee and Yu Jung Lee(Pukyong National University) Multiobjective Fractional Programming with Ratio Invexity

11:10〜11:40 Toshihiko Nishishiraho (University of Ryukyus) Approximation Processes of Bernstein-type Operators

11:40〜12:10 Hidefumi Kawasaki and Jun-ichi Sato (Kyushu University) Discrete Fixed Point Theorems and Their Application to the Game Theory

Lunch Time 13:20〜13:50 Somyot Plubtieng (Naresuan University) Solution of Equilibrium Problems and the Fixed Points of Nonlinear Mappings

13:50〜14:20 Daishi Kuroiwa (Shimane University) On Derivatives of Set-valued Maps in Set Optimization

14:20〜14:50 Satoshi Suzuki and Daishi Kuroiwa (Shimane University) Characterizing Set Containments with Quasiconvex Inequalities

14:50〜15:00 Tea Break 15:00〜15:30 Lai-Jiu Lin and Chih-Sheng Chuang (National Changhua University of Education) Systems of Nonempty Intersection Theorem with Applications

15:30〜16:00 Takamitsu Yamauchi (Shimane University) On the Existence of Continuous Selections Avoiding Extreme Points

16:00〜16:30 Sachiko Atsushiba (Shibaura Institute of Technology) Convergence Theorems for a Family of Relatively Nonexpansive Mappings

16:30〜17:00 Mau-Hsiang Shih (National Taiwan Normal University) Fritz John’s Ellipsoid Theorem: simpler proof and applications September 4th(Tuesday)

9:00〜9:30 Sehie Park (Seoul National University) A Survey on Fixed Point Theorems in Generalized Convex Spaces II

9:30〜10:00 Masao Fukushima (Kyoto University) Semismooth Methods for Linear and Nonlinear Second-Order Cone Programs

10:00〜10:30 Fumiaki Kohsaka (Tokyo Denki University) and Wataru Takahashi (Tokyo Institute of Technology) On Fixed Points of Semicontractive Mappings in Banach Spaces

10:30〜10:40 Tea Break 10:40〜11:10 Suthep Suantai (Chiang Mai University) Common Fixed Points of Finite Family Quasi-nonexpansive Mappings in Banach Spaces

11:10〜11:40 Tetsuzo Tanino (Osaka University) One-Point Solutions Obtained from Best Approximation Problems for Cooperative Games

11:40〜12:10 Saburou Saitoh and Masato Yamada (Gunma University) Nonlinear Transforms and the Theory of Reproducing Kernels

Lunch Time 13:20〜13:50 Jong Kyu Kim and Kyung Soo Kim (Kyungnam University) Convergence Theorems of Implicit Iterative Sequences for a Finite Family of Asymptotically Quasi-Nonexpansive Type Mappings

13:50〜14:20 Yasunori Kimura (Tokyo Institute of Technology) An Equilibrium Problem and Approximation of Its Solutions

14:20〜14:50 Ken-Ichi Mitani (Niigata Institute of Technology), Kichi-Suke Saito (Niigata University) and Tomonari Suzuki (Kyushu Institute of Technology) The James Constant of Absolute Norms on ℝ^{2} II

14:50〜15:00 Tea Break 15:00〜15:30 Shyh Nan Lee(Chun Yuan Christian University) and Mau-Hsiang Shih (National Taiwan Normal University) Octahedral Projection

15:30〜16:00 Tomonari Suzuki (Kyushu Institute of Technology) Meir-Keeler Contraction and L-function

16:00〜16:30 Takayuki Tamura(Chiba University) Mikio Kato (Kyushu Institute of Technology) and Yasuji Takahashi (Okayama Prefectural University) On Infinite Dimensional von Neumann-Jordan Type Constants for Banach Spaces

16:30〜17:00 Jen-Chih Yao (National Sun Yat-sen University) Solution Methods of Variational Inequalities with Pseudomonotone Operators

September 5th(Wednesday)

9:00〜9:30 Sompong Dhompongsa (Chiang Mai University, Thailand) Generalized James Constant and Fixed Point Theorems for Multivalued Nonexpansive Mappings

9:30〜10:00 Hidetoshi Komiya (Keio University) Inequality Comparisons and the Minimax Theorem

10:00〜10:30 Shin-ya Matsushita (Matsue National College of Technology) and Wataru Takahashi (Tokyo Institute of Technology) Existence Theorems for Nonlinear Operators in Banach Spaces

10:30〜10:40 Tea Break 10:40〜11:10 Chin-Tzong Pang (Yuan Ze University) and Mau-Hsiang Shih (National Taiwan Normal University) Simultaneous Schur Stability of Interval Matrices

11:10〜11:40 Koichiro Naito (Kumamoto University) Chaotic Properties of Discrete Dynamical Systems Given by the Gauss Map 11:40〜12:00 Yousuke Araya and Tamaki Tanaka (Niigata University) Existence of Vector Equilibria via Ekeland’s Variational Principle and Its Application

Lunch Time 13:10〜13:40 Attapol Kaewkhao (Burapha University) Some Geometric Conditions and Some Fiexed Point Theorems

13:40〜14:10 Seiji Saito (Doshisha University), Hiroaki Ishii (Osaka University), Kuang-Yih Yeh and Hao-Ching Hsia (National Cheng Kung University) Shortest Paths in Free Spaces Including Obstacles with Fuzzy Boundaries

14:10〜14:40 Yutaka Kimura, Mitsuhiro Hoshino (Akita Prefectural University), Aoi Sato (Miraiya Shoten Co. Ltd.) and Mariko Fujiwara(Akita City Library) A Library Applications by Quantification Analysis

14:40〜15:10 Mitsuhiro Hoshino and Yutaka Kimura (Akita Prefectural University) Quasi-convexity of Model Function and Ordering Process in Self-organizing Maps